Use the table for the question(s) below Day Futures Price 109 107 106 107 104 Suppose oil futures prices are as given in the above table (price per barrel). Suppose you sell 100 crude oil futures contracts, each for 1000 barrels of crude oil, at the current futures price of $108 per barrel on day 0. What is your cumulative profit/loss in your margin account by the end of day 5? O A.-$300,000 B. $300,000 O C. $400,000 O D.-$400,000 O E. SO
Answer #1

Current Future Price per Barrel = $ 108, Number of Barrels per Contract = 1000 and Number of Contracts Sold Short = 100

Total Exposure (in barrels) = 100 x 1000 = 100000

Profit(Loss) at the end of Day 1 = - (Present Day's Price - Previous Day's Price) x Total Exposure =- [(109 - 108) x 100000] = - $ 100000

NOTE: As the speculator has a short position in the futures, he/she loses money if prices appreciate and gain money if prices go down.

Profit(Loss) at the end of Day 2 = - (Present Day's Price - Previous Day's Price) x Total Exposure = -[(107-109) x 100000] = $ 200000

Profit(Loss) at the end of Day 3 = -(Present Day's Price - Previous Day's Price) x Total Exposure = -[(106-107) x 100000] = $ 100000

Profit(Loss) at the end of Day 4 = - (Present Day's Price - Previous Day's Price) x Total Exposure = -[(107-106) x 100000] = - $ 100000

Profit(Loss) at the end of Day 5 = - (Present Day's Price - Previous Day's Price) x Total Exposure = -[(104-107) x 100000] = $ 300000

Cumulative Profit/Loss at the end of Day 5 = -100000 + 200000 + 100000 - 100000 + 300000 = $ 400000

Hence, the correct option is (C)